You bought one of Mastadon Manufacturing Co.’s 6.8 percent coupon bonds one year ago for $1,054. These bonds make annual payments, mature fifteen years from now, and have a par value of $1,000. Suppose you decide to sell your bonds today, when the required return on the bonds is 5 percent.
  
If the inflation rate was 3.6 percent over the past year, what would be your total real return on the investment? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
  
Total real return       _____ %
Answer #1

Bond was purchased one year ago,

Total coupon received = 0.068*1000 = $68

Value of Bond today when YTM = 5%

Maturity of Bond = 15 years from now

Using Bond Calculation,

Price of Bond today = $1,186.83

If bond is sold now,

Total Gain % = 68+(1186.83-1054)/1054

Gain% = 19.05%

As per Fischer Equation,

Real Return = Nominal Return - Inflation Rate

Real Return = 19.05 -3.60

Real Return = 15.45%